OpenStudy (anonymous):

how many numbers between 25 and 400 are divisible by 11? what is their sum?

OpenStudy (anonymous):

who can answer?

OpenStudy (amistre64):

well, it starts with the number 33

OpenStudy (anonymous):

can u pls show the equation sir???

OpenStudy (amistre64):

400-25 = 375 possible results divide that by ll 375/11 = 34.1 so ill assume there is like 34 numbers in that

OpenStudy (anonymous):

and the sum sir?

OpenStudy (amistre64):

there sum can be determined by the expression 33+11n from n=0 to 34 most likely

OpenStudy (anonymous):

wait sir... 33+11n=????

OpenStudy (anonymous):

answer is 7480

OpenStudy (amistre64):

34 11 -- 34 34 ---- 374 11 ---- 385 11 ---- 396 .... found a few more

OpenStudy (amistre64):

pfft, for got to add 33 :) 374 +33 ---- 407 so it looks like 34 is too large and 33 is perfect 0 to 33

OpenStudy (anonymous):

this time i am right i double checked it

OpenStudy (anonymous):

can i ask for the solution sir??? because the solution is greatly needed

OpenStudy (amistre64):

33+11n i s a good expression yes since the first number divisible by 11 is 33

OpenStudy (anonymous):

thanks sambhav

OpenStudy (anonymous):

how to get the n sir???

OpenStudy (amistre64):

\[\sum_{0}^{33}33+11n=\frac{33}{2}(33+374)\]

OpenStudy (anonymous):

a=33 l=396 d=11 n=(396-33/11)+1=35 sum=(35)(2(33)+34(11))/2 =7480 use the formula i gave u earlier

OpenStudy (anonymous):

thanks to to all of you friends :D

OpenStudy (amistre64):

the sum is 7293

OpenStudy (anonymous):

why is sambhav answer is different?

OpenStudy (amistre64):

he most likely has a small error, or i do. but ive got the wolf to dbl chk me :) http://www.wolframalpha.com/input/?i=sum+%2833%2B11n%29%2C+n%3D0..33

OpenStudy (anonymous):

hey @amistre64 ur formu;lla is correctbut ∑03333+11n=332(33+374) =6715.5

OpenStudy (amistre64):

yeah, 374 should have been 396 :) but i accounted for that afterwards

OpenStudy (anonymous):

you used another website for help i used my mind

OpenStudy (amistre64):

at my age, the mind is not so good ;)

OpenStudy (anonymous):

whats the correct answer monsieurs???

OpenStudy (amistre64):

i have to look at my left foot just to remember what color sock i need to put on the right foot :)

OpenStudy (anonymous):

∑03333+11n=(35/2)(33+396)=7507.5

OpenStudy (anonymous):

whats ur age @amistre64

OpenStudy (amistre64):

.... pretty old

OpenStudy (anonymous):

sambhav and amistre what again is the answer???

OpenStudy (amistre64):

7293 is what i get

OpenStudy (anonymous):

pls wait for a minute pls sorry for the argument between us

OpenStudy (anonymous):

its just ok no hard feelings

OpenStudy (amistre64):

0 to 33 is 34 numbers entotal 34(33+396)/2 = 7293

OpenStudy (anonymous):

its right tnx for all your help ive got 2 numbers to go can u cooperate me or help me sir???

OpenStudy (amistre64):

between the both of us, we are sure to work out all the wrinkles ;)

OpenStudy (anonymous):

@bienes_joshua the write answer is 7293

OpenStudy (anonymous):

much thanks of all your help i just need to finish the 2 problems i will close this question and write a new one :D

OpenStudy (amistre64):

starting an index at 1 makes life easier on the count but adds a little bit of complexity to the expression; therefore a modification that is equal is:\[\sum_{1}^{34}33+11(n-1)\]